I have a question that states: find (f of g)(x) and (g of f)(x) for f(x)=x-5 and g(x)=squarerrot x-2. Can anyone help me with this?
Answer by stanbon(48516)

find (f of g)(x) and (g of f)(x) for f(x)=x-5 and g(x)=sqrt(x)-2.
--------
f[g(x)] = f[sqrt(x)-2]
= sqrt(x)-2 -5
= sqrt(x) - 7
------------
g[f(x)]=g[x-5]
=sqrt(x-5) -2
================
Cheers,
Stan H.